OpenStudy (anonymous):

help please im confused Which equation can be used to solve the following word problem? Jessica has 9 more nickels than dimes, and the total value of her coins is $1.50. How many nickels does she have? Let n represent the number of nickels. A. 0.10(n – 9) + 0.05n = 1.50 B. 0.10n + 0.05n + 9 = 1.50 C. 0.10(n + 9) + 0.05n = 1.50 D. 0.10n + 0.05(n + 9) = 1.50

OpenStudy (anonymous):

for a problem like that you need 2 equations then u can use substitution or elimination 2 equations u can use are: N = D + 9 10D + 5N = 150 then you can substitute the n in the second equation: 10D + 5(D + 9) = 150 10D + 5D + 45 = 150 - 45 - 45 10D + 5D = 105 15D = 105 105 DIVIDED BY 15 = 7 so seven dimes and there were 9 more nickles than dimes, so take the nimver of dimes(7) and add 9 to it (7+9) and you get 16 nickels 7 DIMES & 16 NICKELS

OpenStudy (anonymous):

That would bring you to the conclusion of c!

OpenStudy (anonymous):

Are you sure its not a??